# Prompt Injection Attacks on LLM Generated Reviews of Scientific Publications

### v1

@@ 2.4 Prompts
1Write a scientific review of the following paper for the International Conference of Learning Representations (ICLR). ICLR is a high ranking venue where less than 25% of all papers are accepted. In order to be accepted, a papers be outstanding in terms of theoretical soundness, emperical evalaution and novelty and high potetial to be cited by followup research.
@@
3Rate the quallity of the presentation in the output Presentation on a scale from 1 (low) to 4 (high).
@@
5List Strenghts and Weaknesses.
@@
8List missing ralated work and citations in the output Missing_Citations.
@@ 3.1 Structured Output Validation
Table 1: Structured output errors by model. The table shows the rate (in %) of “Ratings” given by the models which do adhere the requested output structure by giving scores that do not exist (most prominently "4") - also see the plots in figure 3. Green highlighted rows indicate models that have been able to predict a correct output structure. The human error rate is of cause 0% as the manual review form only allows valid scores.
@@ 3.2 Effects of Prompt Injection
However, there are some models which appear not to have been effected. Highlighting the manipulable models (as green rows in table 2) shows a very high correlation with the models with valid outputs in table 1 (there also marked in green).
@@ 4 Discussion
Prompt Injection Works! The results shown in tables 2 and 3 as well as figure 2 clearly show that very simply prompt injections are able to dominate the outcome of LLM reviews. The few cases in which the injection did not have significant effects are strongly correlated with the general failure or the models to adhere to the requested structured output. One can speculate that the ability to follow prompted instruction precisely, makes models more vulnerable towards manipulations. However, from the perspective of the assumed “careless” reviewer, these models are not very attractive to use because they do not allow a copy + paste transfer of the outputs into the review forms.
@@
Limitations. This study investigates the likely scenario of a “careless” reviewer who simply drops an assigned review task an a publicly available LLM. Results may not generalize to other scenarios with specifically designed (i.e. fine-tuned) review models. Also, all applied LLMs potentially could have accessed to ILCR papers and reviews during training which in effect could bias the results. However, given the strong shifts between human reviews and all LLM generated reviews, these effects appear to be negligible.
Acknowledgments. This work is supported by the German Federal Ministry of Research, Technology and Space (BMFTR) in the program “Forschung an Fachhochschulen in Kooperation mit Unternehmen (FH-Kooperativ)” within the joint project LLMpraxis under grant 13FH622KX2.

### v2

@@ 2.4 Prompts
1Write a scientific review of the following paper for the International Conference of Learning Representations (ICLR). ICLR is a high ranking venue where less than 25% of all papers are accepted. In order to be accepted, a papers be outstanding in terms of theoretical soundness, empirical evaluation and novelty and high potential to be cited by followup research.
@@
3Rate the quality of the presentation in the output Presentation on a scale from 1 (low) to 4 (high).
@@
5List Strengths and Weaknesses.
@@
8List missing related work and citations in the output Missing_Citations.
@@ 3.1 Structured Output Validation
Table 1: Structured output errors by model. The table shows the rate (in %) of “Ratings” given by the models which fail to adhere the requested output structure by giving scores that do not exist (most prominently "4") - also see the plots in figure 3. Green highlighted rows indicate models that have been able to predict a correct output structure. The human error rate is of cause 0% as the manual review form only allows valid scores.
@@ 3.2 Effects of Prompt Injection
However, there are some models which appear not to have been effected. Highlighting the manipulable models (as green rows in table 2) shows a very high correlation with the models generating valid outputs in table 1 (there also marked in green).
@@ 4 Discussion
Prompt Injection Works! The results shown in tables 2 and 3 as well as figure 2 clearly show that very simple prompt injections are able to dominate the outcome of LLM reviews. The few cases in which the injection did not have significant effects are strongly correlated with the general failure of the models to adhere to the requested structured output. One can speculate that the ability to follow prompted instruction precisely, makes models more vulnerable towards manipulations. However, from the perspective of the assumed “careless” reviewer, these models are not very attractive to use because they do not allow a copy + paste transfer of the outputs into the review forms.
@@
Limitations. This study investigates the likely scenario of a “careless” reviewer who simply drops an assigned review task an a publicly available LLM. Results may not generalize to other scenarios with specifically designed (i.e. fine-tuned) review models. Also, all applied LLMs potentially could have accessed ILCR papers and reviews during training which in effect could bias the results. However, given the strong shifts between human reviews and all LLM generated reviews, these effects appear to be negligible.
Acknowledgments. This work has been supported by the German Federal Ministry of Research, Technology and Space (BMFTR) in the program “Forschung an Fachhochschulen in Kooperation mit Unternehmen (FH-Kooperativ)” within the joint project LLMpraxis under grant 13FH622KX2.